# Whitened Score Diffusion: A Structured Prior for Imaging Inverse Problems

### v3

@@ Appendix C Imaging inverse problems
In our method, the model is trained to denoise correlated noise, and is able to distinguish target image features from anisotropic Gaussian noise features, even though both may share similar spatial frequency support. In contrast, the conventional DM trained to denoise only isotropic Gaussian noise, removes the images features for low enough $\lambda$, as it mistakes the correlated additive noise for the target features. As seen in Fig 6, there are more image features on average outside the additive noise support for CIFAR, and for CelebA, the support of the image features are more closely overlapped with the noise support.
@@
Motion blur is a common image degradation in computational photography. We experiment with a spatially invariant horizontal motion blur kernel of five pixels for CIFAR and seven for CelebA. The additive correlated noise is Gaussian-filtered grayscale WGN with a circular kernel of $\mathrm{std}=2.5$ pixels for CIFAR and 5 pixels for CelebA, each with $\mathrm{SNR}=0.493$. The result is shown in Figs. 5 and 7.
@@
Lens blur is the loss of high spatial frequency information as a result of light rays being focused imperfectly due to the finite aperture size, causing rays from a point source to spread over a region in the image plane rather than converging to a single point. This can be effectively modeled as a convolution between a circular Gaussian kernel and the clean image. In Figs. 5 and 7, we demonstrate our WS diffusion prior on lens deblurring with a Gaussian blur kernel of $\mathrm{STD}=0.8$ and 1.0 for CIFAR and CelebA, respectively. The additive correlated noise is Gaussian-filtered grayscale WGN with a circular kernel of $\mathrm{std}=2.5$ pixels for CIFAR and 5 pixels for CelebA, each with $\mathrm{SNR}=0.810$.
@@
Inverse scattering is a prevalent direction in optical imaging, to recover the permittivity field from measurements under angled illumination. Intensity diffraction tomography (IDT) is a powerful computational microscopy technique that can recover 3D refractive index distribution given a set of 2D measurements. The model can be linearized using the first Born approximation ([28]):
$\mathbf{u}(\mathbf{r})=\mathbf{u}_{i}(\mathbf{r})+\int\mathbf{u}_{i}(\mathbf{r}^{\prime})\mathbf{V}(\mathbf{r}^{\prime})\mathbf{G}(\mathbf{r}-\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime}$ (36)
@@
When the illumination is transmissive, referred to as transmission intensity diffraction tomography (tIDT), meaning that the light passes through the sample, the linear operator, $\mathbf{A}$, results in a mask in the shape of a cross section of a torus that attenuates Fourier coefficients as seen in Figs. 8 and 9, leading to the well-known "missing cone" problem.
In reflection IDT (rIDT), placing the sample object on a specular mirror substrate causes light to reflect towards the camera, enabling the capture of additional axial frequency components and partially filling the missing cone, shown in Figs. 8 and 9.
@@
Differential defocus is a computational imaging technique that aims to recover the depth map from a series of defocused measurements ([2]). The linear operator can be realized with a 2D Laplacian kernel, bandpassing mid-frequency components. In computational microscopy, this is also known as transport of intensity imaging ([49]) to recover the phase and amplitude of an object.
@@ Appendix E Forward Consistency Loss
$L_{2}=\mathbb{E}_{t\sim\mathcal{U}(0,1],\mathbf{x}_{0}\sim p(\mathbf{x}),\mathbf{x}_{t}\sim p(\mathbf{x}_{t}\mid\mathbf{x}_{0})}\left\{\left\|\mathbf{x}_{0}-\frac{\beta_{t}\mathbf{x}_{t}+(1-\alpha_{t}^{2})\mathbf{n}_{\theta}(\mathbf{x}_{t},t)}{\beta_{t}\alpha_{t}}\right\|_{2}^{2}\right\},$ (37)
@@
Empirically, we find that including $L_{2}$ as an auxiliary objective—weighted equally with the primary loss term $L$—leads to improved training stability and faster convergence.

### v4

@@ Appendix C Imaging inverse problems
In our method, the model is trained to denoise correlated noise, and is able to distinguish target image features from anisotropic Gaussian noise features, even though both may share similar spatial frequency support. In contrast, the conventional DM trained to denoise only isotropic Gaussian noise, removes the images features for low enough $\lambda$, as it mistakes the correlated additive noise for the target features. As seen in Fig. 6, there are more image features on average outside the additive noise support for CIFAR, and for CelebA, the support of the image features are more closely overlapped with the noise support. Additional result on the CelebA-HQ ($256\times 256$) dataset is shown in Fig. 7 .
Figure 7: Denoising correlated noise on CelebA-HQ ($256\times 256$). We benchmark our WS DM trained on anisotropic Gaussian noise with the conventional DM (conv) trained on isotropic Gaussian noise. Results for measurements $\mathbf{y}$ with additive grayscale Gaussian noise of $\mathrm{std}=5$ pixels.
@@
Motion blur is a common image degradation in computational photography. We experiment with a spatially invariant horizontal motion blur kernel of five pixels for CIFAR and seven for CelebA. The additive correlated noise is Gaussian-filtered grayscale WGN with a circular kernel of $\mathrm{std}=2.5$ pixels for CIFAR and 5 pixels for CelebA, each with $\mathrm{SNR}=0.493$. The result is shown in Figs. 5 and 8.
@@
Lens blur is the loss of high spatial frequency information as a result of light rays being focused imperfectly due to the finite aperture size, causing rays from a point source to spread over a region in the image plane rather than converging to a single point. This can be effectively modeled as a convolution between a circular Gaussian kernel and the clean image. In Figs. 5 and 8, we demonstrate our WS diffusion prior on lens deblurring with a Gaussian blur kernel of $\mathrm{STD}=0.8$ and 1.0 for CIFAR and CelebA, respectively. The additive correlated noise is Gaussian-filtered grayscale WGN with a circular kernel of $\mathrm{std}=2.5$ pixels for CIFAR and 5 pixels for CelebA, each with $\mathrm{SNR}=0.810$.
@@
Inverse scattering is a prevalent direction in optical imaging, to recover the permittivity field from measurements under angled illumination. Intensity diffraction tomography (IDT) is a powerful computational microscopy technique that can recover 3D refractive index distribution given a set of 2D measurements. The model can be linearized using the first Born approximation ([Ling:18]):
$\mathbf{u}(\mathbf{r})=\mathbf{u}_{i}(\mathbf{r})+\int\mathbf{u}_{i}(\mathbf{r}^{\prime})\mathbf{V}(\mathbf{r}^{\prime})\mathbf{G}(\mathbf{r}-\mathbf{r}^{\prime})\mathrm{d}\mathbf{r}^{\prime},$ (36)
@@
When the illumination is transmissive, referred to as transmission intensity diffraction tomography (tIDT), meaning that the light passes through the sample, the linear operator, $\mathbf{A}$, results in a mask in the shape of a cross section of a torus that attenuates Fourier coefficients as seen in Figs. 9 and 10, leading to the well-known "missing cone" problem.
In reflection IDT (rIDT), placing the sample object on a specular mirror substrate causes light to reflect towards the camera, enabling the capture of additional axial frequency components and partially filling the missing cone, shown in Figs. 9 and 10.
@@
Differential defocus is a computational imaging technique that aims to recover the depth map from a series of defocused measurements ([alexander_focal_2016]). The linear operator can be realized with a 2D Laplacian kernel, bandpassing mid-frequency components. In computational microscopy, this is also known as transport of intensity imaging ([waller_transport_2010]) to recover the phase and amplitude of an object.
@@ Appendix E Forward Consistency Loss
$L_{2}=\mathbb{E}_{t\sim\mathcal{U}(0,1],\mathbf{x}_{0}\sim p(\mathbf{x}),\mathbf{x}_{t}\sim p(\mathbf{x}_{t}\mid\mathbf{x}_{0})}\left\{\left\|\mathbf{x}_{0}-\frac{\beta_{t}\mathbf{x}_{t}+(1-\alpha_{t}^{2})\mathbf{n}_{\theta}(\mathbf{x}_{t},t)}{\beta_{t}\alpha_{t}}\right\|_{2}^{2}\right\}.$ (37)
@@
Empirically, we find that including $L_{2}$ as an auxiliary objective—weighted equally with the primary loss term LL—leads to improved training stability and faster convergence.